OpenStudy (anonymous):

find all of the real 0's and write in factored form. 7x^3-x^2+7x-1

OpenStudy (anonymous):

factor out the x

OpenStudy (anonymous):

i got as far as (7x^3-x2+7x-1)/(7x-1)=x^2+1

OpenStudy (anonymous):

thennnnn....there wouldn't be an answer because they would be imaginary #s :[[!

OpenStudy (anonymous):

is that even right?

OpenStudy (anonymous):

\[7 x^3-x^2+7 x-1=(7 x-1) \left(x^2+1\right) \]A plot is attached.